OpenStudy (anonymous):

Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. y = x^2 + 2x + 1 and y = x^2 + 3x + 2

OpenStudy (mertsj):

y=(x+1)(x+1) y=(x+2)(x+1)

OpenStudy (anonymous):

teach me how you got that? :)

OpenStudy (mertsj):

x^2+2x+1 is a trinomial. We know that trinomals come from multiplying two binomials. Are you with me so far?

OpenStudy (anonymous):

yuup

OpenStudy (mertsj):

Do you know how to multiply two binomials?

OpenStudy (anonymous):

no... please do explain haha read my username :P

OpenStudy (mertsj):

Have you heard of FOIL?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

You multiply two binomials by using FOIL which means you multiply the First terms of the two binomials. O stands for Outer which means to multiply the outer terms of the two binomials. I stands for Inner which means you multiply the inner terms of the two binomials. L stands for Last which means you multiply the Last terms of the two binomials.

OpenStudy (anonymous):

okay... so what im a foiling?

OpenStudy (mertsj):

So when you factor a trinomial, you reverse that procedure. You have to find the two binomials that were multiplied to get the trinomial. But if you know how to FOIL, then you can figure out what the two binomials are.

OpenStudy (mertsj):

For example: If I want to find two binomials that will multiply together to get x^2+2x+1 I know they both have to start with x: (x )(x )

OpenStudy (mertsj):

Because x times x is x^2

OpenStudy (anonymous):

ok im following

OpenStudy (mertsj):

If I want to get the 1 on the end, then it would have to be 1 times 1 or -1 times -1

OpenStudy (mertsj):

(x+1)(x+1) or (x-1)(x-1)

OpenStudy (mertsj):

Then I check out the Outer and Inner to see which one gives me +2x.

OpenStudy (mertsj):

Clearly it is the (x+1)(x+1)

OpenStudy (anonymous):

makes sense

OpenStudy (mertsj):

So those are the factors of x^2+2x+1 because when I multiply (x+1)(x+1) using FOIL, I get x^2+2x+1

OpenStudy (mertsj):

So if I wanted to factor x^2+3x+2 once again the two binomials would have to start with x: (x )(x )

OpenStudy (mertsj):

How could I get +2 for the last term?

OpenStudy (anonymous):

x + 2

OpenStudy (mertsj):

I could get +2 for the last term by multiplying 2 times 1 or -2 times -1 Which choice should I make and why?

OpenStudy (anonymous):

2 times 1 will get you positive 2

OpenStudy (mertsj):

So will -2 times -1

OpenStudy (anonymous):

so why would either be wrong?

OpenStudy (mertsj):

Because only one of them will result in +3x for the Outer + the Inner

OpenStudy (anonymous):

so you need the positive ones right?

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (mertsj):

the first one has two distinct roots. The second one had 1 double root.

OpenStudy (anonymous):

thanks so much!!!